### v2

@@ Abstract
We investigate the distribution of subhalos in the vicinity of dark matter halo pairs in the Illustris-TNG simulation. The satellites around halo pairs with separations $d_{\text{sep}}>0.5~{}h^{-1}\text{Mpc}$ exhibit a non-uniform distribution, which is a combined result of two distinct features: the “bulging” distribution, characterized by an overabundance along the pairwise direction, and the “lopsided” distribution, which shows an overabundance in the region between the paired halos. The bulging signal is stronger for halo pairs that are more widely separated and involve less massive halos, while the lopsidedness strengthens as the halo mass increases. Both signals depend weakly on halo mass ratio, and are primarily contributed by subhalos that are relatively distant from host halos. These measurements can be broadly reproduced by the overlap effect, provided the spatial alignment of halos is properly taken into account. Our results imply that the lopsidedness and bulging distribution may arise from two distinct origins: the former is caused simply by the overlap of two closely placed halos, while the latter is driven by the alignment of halos with large-scale filamentary structures. We examine the impact of projection and selection effects by conducting the same analysis of satellite distribution in two dimensions and in a mock catalog that replicates the selection effects of the SDSS galaxy sample. We find that the 3D-to-2D projection significantly suppresses the bulging distribution, with particularly strong effects at large pair separations, low halo masses, and large halo mass ratios.
@@ 1 Introduction
In a parallel paper (Guo et al. 2025; hereafter Paper I), we have extended the study of lopsided and bulging distributions of satellites around halo pairs, by selecting a sample of paired central galaxies from the Sloan Digitak Sky Survey (SDSS; York et al., 2000) and examining the lopsidedness and bulging signals as function of host halo mass, halo mass ratio and pair separation. In addition, we have constructed a mock catalog that has the same selection effects as the observational sample, based on the Illustris-TNG300 simulation. By incorporating the orientation of individual galaxies and the alignment of galaxy/hao orientation with halo pair orientation, our mock catalog successfully reproduces the lopsided and bulging distributions of satellites across halo pairs of various properties. Our results demonstrate that the angular distribution of satellites is a natural phenomenon of the $\Lambda$CDM model, eliminating the need to introduce other mechanisms such as the local gravitational effects proposed by Libeskind et al. (2016).
@@ 2.1 Satellite distribution around halo pairs
Figure 1 presents a schematic diagram of the halo pairs. Following Gong et al. (2019), we consider all the subhalos located within $d_{\mathrm{sep}}/2$ from the nearer halo when measuring the satellite distribution for a halo pair. This includes not only the subhalos within the virial radii of the two halos but also some of those located outside (as indicated by the blue and red colors in the figure). For each subhalo in consideration, we measure its positional angle $\theta$, defined as the angle between its position vector with respect to the nearer halo and the line connecting the two halos. Defined this way, angles in the range of $0\leq\theta\leq 90^{\circ}$ (or equivalently $0\leq\cos\theta\leq 1$) correspond to subhalos located in the internal region between the two halos, while angles in the range of $90^{\circ}<\theta\leq 180^{\circ}$ (or equivalently $-1\leq\cos\theta<0$) correspond to those located on the outer sides.
Figure 2: Satellite distribution in 3D as measured by $P(\cos\theta)$ is shown for halo pairs of varying separation dsep, primary halo mass Mp, and secondary-to-primary halo mass ratio Ms/Mp, as indicated. The horizontal line indicates the uniform distribution.
Figure 3: Satellite distribution $P(\cos\theta)$ is shown for halo pairs of different primary halo masses, as well as different satellite-to-host distance for given halo mass, as indicated.
Figure 2 displays the distribution of $\cos\theta$ for halo pairs with varying primary halo masses ($M_{p}$), pair separations ($d_{\text{sep}}$), and secondary-to-primary halo mass ratios ($M_{s}/M_{p}$). Additionally, Figure 3 compares the $\cos\theta$ distribution for subhalos at different distances from their host halo, as measured for halo pairs in different primary halo mass bins. Overall, $\cos\theta$ presents a non-uniform distribution for halo pairs with separations $d_{\text{sep}}>0.5~{}h^{-1}\text{Mpc}$, showing a minimum of $P(\cos\theta)$ around $\cos\theta=0$ and increased values towards both $\cos\theta=1$ and $\cos\theta=-1$. This indicates a “bulging” distribution of satellites around the halo pairs, characterized by an overabundance of satellites along the connection line of the halo pairs, both inside and outside the pairs. The bulging signal is stronger for halo pairs that are more widely separated and involve less massive halos, but it exhibits a rather weak dependence on the secondary-to-primary halo mass ratio. As shown by Figure 3, the bulging signal at a given halo mass is primarily contributed by subhalos that are relatively distant from their host halos.
@@ 2.2 Satellite distribution around overlapping halos
One possible explanation for the lopsided and bulging distribution of satellites around paired halos is the so-called “overlap effect”: an overabundance of satellites is naturally found between two halos when they are brought closer together, along with their surrounding satellite distribution. In previous studies, this hypothesis has been tested by comparing the satellite distribution around actual halo pairs with that around overlapping halos. For each actual halo pair, two halos with masses similar to those of the halos in the pair are selected from the simulation, and along with their surroundings, they are placed at the same separation as the actual halo to form the corresponding overlapping halos. In practice, we require the overlapping halos to be closely matched in mass with the corresponding halos in the actual pairs, with a tolerance of $\Delta\log_{10}M_{h}\leq 0.1$, and to be separated at the same distance as the actual halo pair, with a tolerance of $\Delta d_{\text{sep}}\leq 0.5~{}h^{-1}\text{Mpc}$. For each of the actual pairs, this process is repeated 20 times, resulting in a sample of overlapping halos that is 20 times the size of the actual pair sample.
@@
To enable quantitative comparisons, we define two parameters, $P$ and $Q$, to respectively quantify the strength of the bulging and lopsidedness signals. Specifically, $P$ is defined as the integral of $P(\cos\theta)$ over the ranges $1.00.5<\cos\theta<$ 1.0 and $-1.0<\cos\theta<-0.5$, divided by the integral of $P(\cos\theta)$ over the range $-0.5<\cos\theta<0.5$. This parameter measures the relative abundance of satellites along the direction of the pair orientation compared to that perpendicular to it. Following Gong et al. (2019), the $Q$ parameter is defined as the ratio of $P(\cos\theta)$ at $1.00.8<\cos\theta<$ 1.0 to that at $-1.0<\cos\theta<-0.8$, thereby measuring the relative satellite abundance between the inner and outer sides of the halo pair. The $P$ and $Q$ parameters are calculated for all the $P(\cos\theta)$ measurements, as indicated in each panel of Figure 4, Figure 5, and Figure 6.
@@ 3.1 Satellite distribution around halo pairs
For the 2D analysis, we take the $x-y$ plane of the IllustrisTNG-ODM simulation as the projection plane. We have verified that our results remain unchanged when using the $x-z$ or $y-z$ planes. We select halo pairs on the 2D plane, requiring them to have halo masses in the same range as those analyzed in 3D. Additionally, these pairs must have projected separations in the range of $0<d^{\mathrm{proj}}_{\mathrm{sep}}<2~{}h^{-1}\mathrm{Mpc}$ and velocity differences $d_{\text{v}}<1000~{}\mathrm{km/s}$ along the line of sight (LoS), specifically the $z$-axis. We account for both the position difference and the peculiar velocity difference when calculating the velocity difference $d_{\text{v}}$ between two given halos. Similar to the halo pairs selected in 3D, each of the projected halo pairs is additionally required to have no third halo within a cylinder centered at the midpoint of the halo pair, which has a projected radius equal to $d^{\mathrm{proj}}_{\mathrm{sep}}$ and a LoS depth of ±$\pm 1000~{}\mathrm{km/s}$. All subhalos located within $0.5d^{\mathrm{proj}}_{\mathrm{sep}}$ in projection and ±$\pm 500~{}\mathrm{km/s}$ along LoS from the nearer halo are considered satellites around the halo pair. For each satellite, we calculate the angle in the projection plane, $\Theta$, defined as the angle between the line connecting the satellite to the nearer host halo and the line connecting the two halos. Similar to the case in 3D, a satellite is labeled as being located “inside” the halo pair if $\Theta\in(0^{\circ},90^{\circ})$ and as being “outside” if $\Theta\in(90.0^{\circ},180.0^{\circ})$.
@@
[…]
@@ 4.2 Implications for the origin of bulging and lopsidedness
Our comparison between the 3D and 2D analyses demonstrates that, although the projection effect weakens the bulging distribution by a factor that depends on pair separation and halo mass, the bulging and lopsided distributions of satellites are fundamentally intrinsic behaviors of the halo pairs in a $\Lambda$CDM universe. More importantly, we find that the satellite distribution around halo pairs with varying properties can be reasonably well reproduced by the overlap effect, either with or without including the spatial alignment of halos. Specifically, the overlapping halos that do not consider halo alignment cannot produce any bulging signal, but they successfully replicate the satellite distribution for halo pairs that are closely separated or involve massive halos. In contrast, the overlapping halos that are aligned with each other can produce significant bulging distributions that match the satellite distribution around halo pairs with relatively large separations and low halo masses.
@@
The excess of matter between dark matter halo pairs has been well established in both simulations and observations, and this excess has indeed been utilized as a common method to identify filaments (Colberg et al., 2005; Clampitt et al., 2016; Epps & Hudson, 2017; Xia et al., 2020; Kondo et al., 2020; Yang et al., 2022). Additionally, previous studies have found that galaxy pairs (Tempel & Tamm, 2015; Mesa et al., 2018; Sarkar & Pandey, 2024) and galaxy multiplets composed of small groups of 2-4 member galaxies (Rong et al., 2024; Lamman et al., 2024) tend to align with their host filaments, with stronger alignment observed for more widely separated halo pairs (e.g. Tempel & Tamm, 2015). Moreover, the thickness of filaments is found to be approximately 1-2 Mpc in the local universe (e.g. Aragón-Calvo et al., 2010; Bond et al., 2010; Cautun et al., 2014; Wang et al., 2024), which is consistent with our finding that the bulging signal dominates over the lopsidedness for wide pairs with $d_{\text{sep}}\gtrsim 1~{}h^{-1}\text{Mpc}$. All of these results align well with the idea that the bulging distribution has an alignment origin, as described above. This picture is also consistent with our finding that the bulging distribution is primarily contributed by subhalos that are relatively distant from their host halos, considering that the satellites in the outskirts of halos are more recently accreted and thus more capable of keeping their anisotropic distribution inherited from the filaments (Li et al. 2013; Gong et al. 2019; Liu et al. 2024).
Furthermore, although the alignment of isolated halos has previously been found to be positively correlated with halo mass (e.g. Faltenbacher et al., 2008), we find that the bulging of satellites becomes more pronounced in halo pairs of lower masses at fixed pair separation. This implies that pairs of halos with relatively low masses are better tracers of the large-scale filamentary structures compared to those involving massive halos, which are preferentially located in high-density knots of the cosmic web that connect multiple filaments (e.g. Aragón-Calvo et al., 2010; Codis et al., 2018; Gouin et al., 2021).
@@ 4.3 Projection and selection effect
In addition, the 3D-to-2D projection affects the satellite distribution around actual halo pairs and that around overlapping halos in different ways. For instance, as can be seen from Figure 9, although the overlap sample and the actual pair sample agree well with each other when analyzed in 3D, their 2D measurements obtained with the mock catalog present significant discrepancies, particularly for widely-separated pairs of low-mass halos. This echos the similar discrepancies found between halo pairs and overlapping halos from the SDSS galaxy sample; see Fig. 11 and discussion at the end of Section 4 in Paper I.

### v3

@@ Abstract
We use the Illustris-TNG simulation to investigate the anisotropic distribution of subhalos in/around dark matter halo pairs. We measure the position angle ($\theta$) of each subhalo by the angle between the line connecting it to the nearest host halo and the line connecting the paired halos, and examine $P(\cos\theta)$ (the distribution of $\cos\theta$ of all subhalos) for halo pairs with various separations ($d_{\text{sep}}$), primary halo masses ($M_{p}$) and secondary-to-primary halo mass ratios ($M_{s}/M_{p}$). We find that $P(\cos\theta)$ generally exhibits a combined result of two distinct features: the “bulging” distribution characterized by an overabundance along the pairwise direction, and the “lopsided” distribution showing an overabundance in the region between the paired halos. The bulging signal is stronger for halo pairs with larger $d_{\text{sep}}$ and smaller $M_{p}$, while the lopsidedness strengthens as $M_{p}$ increases. Both signals depend weakly on $M_{s}/M_{p}$, and are primarily contributed by subhalos that are relatively distant from host halos. Remarkably, these measurements can be broadly reproduced by the overlap effect, provided the spatial alignment of halos is properly taken into account. Our findings suggest distinct origins: lopsidedness arises from simple halo overlap, while bulging reflects alignment with large-scale filaments. We examine the impact of projection and selection effects by conducting the same analysis in two dimensions and in a mock catalog that replicates the selection effects of the SDSS galaxy sample. We find that the 3D-to-2D projection significantly suppresses the bulging distribution, with particularly strong effects at large $d_{\text{sep}}$, small $M_{p}$, and large $M_{s}/M_{p}$.
@@ 1 Introduction
[…]
@@ 2.1 Satellite distribution around halo pairs
Figure 1 presents a schematic diagram of the halo pairs. Following Gong et al. (2019), we consider all the subhalos located within $d_{\mathrm{sep}}/2$ from the nearer halo when measuring the satellite distribution for a halo pair. For the range of $d_{\mathrm{sep}}$ considered, we establish that $d_{\mathrm{sep}}/2>R_{p,\mathrm{vir}}$ (the virial radius of the primary halo) holds for nearly all halo pairs. This criterion ensures the representative inclusion of subhalos residing within the virial radii of both halos. Furthermore, our satellite definition encompasses subhalos located beyond these virial radii, as visually indicated by the blue and red ellipses in the figure. These external subhalos are significant as they probe the connection between the halo pairs and the surrounding large-scale structure. For each subhalo in consideration, we measure its positional angle $\theta$, defined as the angle between its position vector with respect to the nearer halo and the line connecting the two halos. Defined this way, angles in the range of $0\leq\theta\leq 90^{\circ}$ (or equivalently $0\leq\cos\theta\leq 1$) correspond to subhalos located in the internal region between the two halos, while angles in the range of $90^{\circ}<\theta\leq 180^{\circ}$ (or equivalently $1\leq\cos\theta<0$) correspond to those located on the outer sides.
Figure 2: Satellite distribution in 3D as measured by $P(\cos\theta)$ is shown for halo pairs of varying separation dsep, primary halo mass Mp, and secondary-to-primary halo mass ratio Ms/Mp, as indicated. The horizontal line indicates the uniform distribution. The numbers listed at the bottom are the sample size of the corresponding subsamples.
Figure 3: Satellite distribution $P(\cos\theta)$ is shown for halo pairs of different primary halo masses, as well as different satellite-to-host distance for given halo mass, as indicated. Those two subsamples have the same number of pairs, as listed at the bottom.
Figure 2 displays the distribution of $\cos\theta$ for halo pairs with varying primary halo masses ($M_{p}$), pair separations ($d_{\text{sep}}$), and secondary-to-primary halo mass ratios ($M_{s}/M_{p}$). Additionally, Figure 3 compares the $\cos\theta$ distribution for subhalos at different distances from their host halo, as measured for halo pairs in different primary halo mass bins. In both figures and throughout the paper, errors of $P(\cos\theta)$ are estimated with the technique of bootstrap resampling, by randomly resampling 50 times the actual halo pairs or the overlapping halos in a given sample. Overall, $\cos\theta$ presents a non-uniform distribution for halo pairs with separations $d_{\text{sep}}>0.5~h^{-1}\text{Mpc}$, showing a minimum of $P(\cos\theta)$ around $\cos\theta=0$ and increased values towards both $\cos\theta=1$ and $\cos\theta=-1$. This indicates a “bulging” distribution of satellites around the halo pairs, characterized by an overabundance of satellites along the connection line of the halo pairs, both inside and outside the pairs. The bulging signal is stronger for halo pairs that are more widely separated and involve less massive halos, but it exhibits a rather weak dependence on the secondary-to-primary halo mass ratio. As shown by Figure 3, the bulging signal at a given halo mass is primarily contributed by subhalos that are relatively distant from their host halos.
@@
It might be questioned whether the relatively wide ranges of $d_{\text{sep}}$, $M_{p}$, and $M_{s}/M_{p}$ have introduced biases into our results, given the potential dependence of $d_{\text{sep}}$ on $M_{p}$ and $M_{s}/M_{p}$. To address this, we have examined the distribution of halo pairs in our sample on the $d_{\text{sep}}$ versus $M_{p}$ and $d_{\text{sep}}$ versus $M_{s}/M_{p}$ planes. Overall, $d_{\text{sep}}$ exhibits weak positive correlations with $M_{p}$ and weak negative correlations with $M_{s}/M_{p}$. However, when restricted to individual subsamples, these correlations become considerably weaker, indicating a negligible influence on the results presented in this paper.
@@ 2.2 Satellite distribution around overlapping halos
One possible explanation for the lopsided and bulging distribution of satellites around paired halos is the so-called “overlap effect”: an overabundance of satellites is naturally found between two halos when they are brought closer together, along with their surrounding satellite distribution. In previous studies, this hypothesis has been tested by comparing the satellite distribution around actual halo pairs with that around overlapping halos (Libeskind et al. 2016; Pawlowski et al. 2017; Gong et al. 2019). For each actual halo pair, two halos with masses similar to those of the halos in the pair are selected from the simulation, and along with their surroundings, they are placed at the same separation as the actual halo to form the corresponding overlapping halos. In practice, we require the overlapping halos to be closely matched in mass with the corresponding halos in the actual pairs, with a tolerance of $\Delta\log_{10}M_{h}\leq 0.1$, and to be separated at the same distance as the actual halo pair, with a tolerance of $\Delta d_{\text{sep}}\leq 0.5~h^{-1}\text{Mpc}$. For each of the actual pairs, this process is repeated 200 times, resulting in a sample of overlapping halos that is 200 times the size of the actual pair sample.
@@
To enable quantitative comparisons, we define two parameters, $P$ and $Q$, to respectively quantify the strength of the bulging and lopsidedness signals. Specifically, $P$ is defined as the integral of $P(\cos\theta)$ over the ranges $0.5<\cos\theta<1.0$ and $1.0<\cos\theta<-0.5$, divided by the integral of $P(\cos\theta)$ over the range $0.5<\cos\theta<0.5$. This parameter measures the relative abundance of satellites along the direction of the pair orientation compared to that perpendicular to it. Following Gong et al. (2019), the $Q$ parameter is defined as the ratio of $P(\cos\theta)$ at $0.8<\cos\theta<1.0$ to that at $1.0<\cos\theta<-0.8$, thereby measuring the relative satellite abundance between the inner and outer sides of the halo pair. The $P$ and $Q$ parameters are calculated for all the $P(\cos\theta)$ measurements, as indicated in each panel of Figure 4, Figure 5, and Figure 6.
@@ 3.1 Satellite distribution around halo pairs
For the 2D analysis, we take the xyx-y plane of the IllustrisTNG-ODM simulation as the projection plane. We have verified that our results remain unchanged when using the xzx-z or yzy-z planes. We select halo pairs on the 2D plane, requiring them to have halo masses in the same range as those analyzed in 3D. Additionally, these pairs must have projected separations in the range of $0<d^{\mathrm{proj}}_{\mathrm{sep}}<2~h^{-1}\mathrm{Mpc}$ and velocity differences $d_{\text{v}}<1000~\mathrm{km/s}$ along the line of sight (LoS), specifically the $z$-axis. We account for both the position difference and the peculiar velocity difference when calculating the velocity difference $d_{\text{v}}$ between two given halos. Similar to the halo pairs selected in 3D, each of the projected halo pairs is additionally required to have no third halo within a cylinder centered at the midpoint of the halo pair, which has a projected radius equal to $d^{\mathrm{proj}}_{\mathrm{sep}}$ and a LoS depth of $\pm 1000~\mathrm{km/s}$. All subhalos located within $0.5d^{\mathrm{proj}}_{\mathrm{sep}}$ in projection and $\pm 500~\mathrm{km/s}$ along LoS from the nearer halo are considered satellites around the halo pair. For each satellite, we calculate the angle in the projection plane, $\Theta$, defined as the angle between the line connecting the satellite to the nearer host halo and the line connecting the two halos. Similar to the case in 3D, a satellite is labeled as being located “inside” the halo pair if $\Theta\in(0^{\circ},90^{\circ})$ and as being “outside” if $\Theta\in(90.0^{\circ},180.0^{\circ})$.
@@
In Figure 7 and Figure 8, we compare the satellite distribution around halo pairs as measured by $P(\Theta)$ on the projection plane (red lines) and in the mock catalog (blue symbols), for halo pairs with different separations, halo masses, and halo mass ratios. Note that the two lowest mass bins from previous figures have been combined into a single mass bin, $11.6<\mathrm{log}{10}(M{\mathrm{p}}[\mathrm{M}_{\odot}/h])<12.6$, in order to avoid the relatively large Poisson noise present in those narrow bins in the mock catalog. As can be seen, the satellites selected in 2D and those from the mock catalog display very similar distributions around halo pairs with relatively small separations ($d_{\text{sep}}^{\text{proj}}<1~h^{-1}\text{Mpc}$). At larger separations, however, the satellite distributions in 2D are slightly less bulging and lopsided than those in the mock catalog, with the discrepancy appearing to increase as $d_{\text{sep}}^{\text{proj}}$ increases. The relatively small differences found between the measurements from 2D and the mock catalog indicate that the 3D-to-2D projection is the dominant effect in a redshift survey like SDSS, and selection effects other than projection should not introduce significant biases, provided they are properly accounted for in the statistical measurement, as we have done in both Paper I and this work.
@@ 4.2 Implications for the origin of bulging and lopsidedness
[…]
@@
The excess of matter between dark matter halo pairs has been well established in both simulations and observations, and this excess has indeed been utilized as a common method to identify filaments (Colberg et al., 2005; Clampitt et al., 2016; Epps & Hudson, 2017; Xia et al., 2020; Kondo et al., 2020; Yang et al., 2022). Additionally, previous studies have found that galaxy pairs (Tempel & Tamm, 2015; Mesa et al., 2018; Sarkar & Pandey, 2024) and galaxy multiplets composed of small groups of 2-4 member galaxies (Rong et al., 2024; Lamman et al., 2024) tend to align with their host filaments, with stronger alignment observed for more widely separated halo pairs (e.g. Tempel & Tamm, 2015). Moreover, the thickness of filaments is found to be approximately 1-2 Mpc in the local universe (e.g. Aragón-Calvo et al., 2010; Bond et al., 2010; Cautun et al., 2014; Wang et al., 2024), which is consistent with our finding that the bulging signal dominates over the lopsidedness for wide pairs with $d_{\text{sep}}\gtrsim 1~h^{-1}\text{Mpc}$. All of these results align well with the idea that the bulging distribution has an alignment origin, as described above. This picture is also consistent with our finding that the bulging distribution is primarily contributed by subhalos that are relatively distant from their host halos, considering that the satellites in the outskirts of halos are more recently accreted and thus more capable of keeping their anisotropic distribution inherited from the filaments (Li et al. 2013; Gong et al. 2019).
One may wonder whether and how the observed lopsided and bulging satellite distributions around halo pairs relate to those found around isolated halos/galaxies. Previous studies attribute both anisotropic distributions around isolated systems to large-scale tidal fields and anisotropic accretion (references in section 1). Our analysis reveals a distinct physical origin for lopsidedness in paired systems: it arises primarily from geometric overlap effects due to halo alignment, rather than contributions from isolated-halo lopsidedness or local gravitational mechanisms. This contrasts with the anisotropic accretion-driven lopsidedness found in isolated galaxies (e.g. Liu et al., 2024). For the bulging distribution, we find that satellite alignment with large-scale filaments—likely tied to the orientation of paired halos—provides a plausible explanation. This mechanism resembles the origin of bulging distributions around isolated halos, whose orientations also tend to align with the large-scale dark matter distribution (e.g. Faltenbacher et al., 2008). However, a key difference emerges in the mass dependence: while isolated halo alignment strengthens with increasing halo mass, the bulging signal in paired systems intensifies for lower-mass pairs at fixed separation. This inverse mass trend has two implications. First, low-mass halo pairs trace large-scale filaments more effectively than massive pairs, which preferentially reside in high-density knots connecting multiple filaments (e.g. Aragón-Calvo et al., 2010; Codis et al., 2018; Gouin et al., 2021). Second, the contrasting mass dependencies may indicate distinct physical origins for bulging in paired versus isolated systems. While our model of aligned overlapping halos broadly explains the satellite distribution around halo pairs of various properties, discrepancies remain unresolved in certain cases. A direct, systematic comparison of satellite anisotropy between isolated and paired systems would be essential to fully understand their distribution mechanisms in both contexts.
@@ 4.3 Projection and selection effect
In addition, the 3D-to-2D projection affects the satellite distribution around actual halo pairs and that around overlapping halos in different ways. The projection effects exert a significantly stronger influence on overlapping systems compared to the pairs. This disparity primarily arises because the projection distance exceeds the intrinsic 3D separation. On the one hand, the bulging signal is stronger for the wider-separated pair system, thereby eliminating some projected effects from the interiors of the satellites. On the other hand, the bulging signal is further weakened for overlaps due to the weaker bulging feature for larger separation of the overlaps. For instance, as can be seen from Figure 9, although the overlap sample and the actual pair sample agree well with each other when analyzed in 3D, their 2D measurements obtained with the mock catalog present significant discrepancies, particularly for widely-separated pairs of low-mass halos. This echos the similar discrepancies found between halo pairs and overlapping halos from the SDSS galaxy sample; see Fig. 11 and discussion at the end of Section 4 in Paper I.